OpenStudy (anonymous):

How do I find the indefinite integral of: (3x-1)/(x^2-x+1)? Thanks.

OpenStudy (anonymous):

\[\int\limits_{}\frac{ 3x-1 }{ x^2-x+1 }dx{}\]

OpenStudy (anonymous):

You're gonna wanna "decompose" that fraction in the integrand. \[ \frac{ ax+b }{ (x-r)(x-t) } = \frac{ c }{ (x-r) } + \frac{ d }{ (x-t) }\] where a,b,r,t are known numbers from the original fraction; and c,d are unknowns. Working through the algebra you get (the dx and dr are not differentials!) \[cx+ct+dx+dr = ax+b\] This only holds up when all the x terms are equal and all the constant terms are equal, so we can break this into 2 equations. \[c + d = a\] \[ct + dr = b\] Some algebra later and you get \[d = \frac{ b-at }{ r-t }\] \[c=\frac{ ar-b }{ r-t }\] That should help you simplify your integrand. From there its just a straightforward u sub to evaluate the integral. Let me know if you need anymore help.

OpenStudy (anonymous):

Partial fraction decomp may be a bit tricky here, as there are no real roots to the denominator's polynomial. I would suggest the following: \[\int\frac{3x-1}{x^2-x+1}~dx=\int\frac{x}{x^2-x+1}~dx+\int\frac{2x-1}{x^2-x+1}~dx\] A simple \(u\)-sub works for the last integral: \(u=x^2-x+1\), so \(du=(2x-1)~dx\). For the first integral, you would complete the square in the denominator: \[\begin{align*}x^2-x+1&=x^2-x+\frac{1}{4}+\frac{3}{4}\\\\ &=\left(x-\frac{1}{2}\right)^2+\frac{3}{4} \end{align*}\] Next step involves a preliminary algebraic sub: \(x-\dfrac{1}{2}=\dfrac{\sqrt3}{2}\tan t\), so that \(dx=\dfrac{\sqrt3}{2}\sec^2t~dt\). \[\begin{align*}\int\frac{x}{\left(x-\frac{1}{2}\right)^2+\frac{3}{4}}~dx&=\int\frac{\frac{\sqrt3}{2}\tan t+\frac{1}{2}}{\left(\frac{\sqrt3}{2}\tan t\right)^2+\frac{3}{4}}\times\dfrac{\sqrt3}{2}\sec^2t~dt\\ &=\int\frac{\frac{\sqrt3}{2}\tan t+\frac{1}{2}}{\frac{3}{4}\tan^2t+\frac{3}{4}}\times\dfrac{\sqrt3}{2}\sec^2t~dt\\ &=\frac{\frac{\sqrt3}{2}}{\frac{3}{4}}\int\frac{\frac{\sqrt3}{2}\tan t+\frac{1}{2}}{\tan^2t+1}\times\sec^2t~dt\\ &=\frac{2\sqrt3}{3}\int\frac{\frac{\sqrt3}{2}\tan t+\frac{1}{2}}{\sec^2t}\times\sec^2t~dt\\ &=\frac{2\sqrt3}{3}\int\left(\frac{\sqrt3}{2}\tan t+\frac{1}{2}\right)~dt\\ &=\int\tan t~dt+\frac{1}{2}\int dt\\\\\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

Thanks! Any reason I cant see the equations? all i see is code ("frac\"..)

OpenStudy (anonymous):

Likely a browser problem? Try switching your browser, or checking back later. I'm sure you can understand why I'd be reluctant to recreate that as a drawing haha